Marga (35 kg) and Kasey (45kg) were playing catch. They agree to toss a 0.50kg ball towards each other. If the ball is moving at a rate of 0.50m/s, how fast must Marga and Kasey move in order to catch the ball? Type in your complete solution. Clue: The ball does not move once it was caught. *


1
Expert's answer
2021-06-28T16:48:06-0400
"m_Mv_M=2mv\\35v_M=2(0.5)(0.5)\\v_M=0.014\frac{m}{s}"

"m_Kv_K=2mv\\45v_K=2(0.5)(0.5)\\v_K=0.011\frac{m}{s}"